A 12-by-12 in. open pan containing water at 140 °F is placed in the exact center of a 16-by-20 ft room having a ceiling temperature of 75°F. If the distance from the water surface to the ceiling is 10 ft and the ceiling is painted with an oil paint, what is the net radiant-heat exchange between the water surface and the ceiling?
1
Expert's answer
2021-10-30T11:20:08-0400

"T_1=140"

"T_2=75"

"\epsilon_1=0.9"

"\epsilon_2=0.4"

"\sigma=5.67*10^{-8}"

"Q_{net}=\frac{\sigma(T_1^4-T_2^4)}{\frac{1}{\epsilon_1}+\frac{1}{\epsilon_2}-1}"

"=\frac{5.67*10^{-8}(140^4-75^4)}{\frac{1}{0.9}+\frac{1}{0.4}-1}"

"=7.6549BTU\/hrft"